Circular No. 6846

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

POSSIBLE NOVA IN SAGITTARIUS W. Liller, Viña del Mar, Chile, reports his discovery of an apparent

nova (mag 7.8) on Technical Pan photographs taken with PROBLICOM on Mar. 22.3628, 22.3642, and 22.3663 UT with an 85-mm camera lens (+ orange filter). The star is located at α = 18^{h}21^{m}.6, δ = −27^{o}32′ (equinox 2000.0). Nothing brighter than mag 11.5 appears at this location on two photographs taken on Mar. 18.35. On Mar. 23.316, the CCD magnitude was measured by Liller as V = 7.74. A low-resolution spectrogram obtained with a 0.20-m Schmidt camera (+ objective prism + CCD) on Mar. 23 shows Hα in emission at a level ∼ 1.5 times brighter than the surrounding continuum.

SAX J1747.0–2853 AND GX +0.2–0.2 J. in ’t Zand, Space Research Organization of the Netherlands (SRON),

Utrecht; A. Bazzano, M. Cocchi, and P. Ubertini, Istituto di Astrofisica Spaziale, CNR, Rome; J. M. Muller, BeppoSAX Science Data Center, Rome, and SRON; and V. Torroni, BeppoSAX Science Operation Center, Rome, communicate: “A transient, bursting x-ray source was detected in an observation of the Galactic bulge with the Wide Field Camera 2 onboard BeppoSAX between Mar. 20.3 and 22.1 UT. This object, designated SAX J1747.0–2853, is at α = 17^{h}47^{m}02^{s}, δ = −28^{o}53′.0 (equinox 2000.0; preliminary error radius 3′), and the position is 0′.3 from the centroid position of GX +0.2–0.2, a transient last seen in 1976 (Proctor et al. 1978, MNRAS 185, 745). This supports the identification of SAX J1747.0–2853 with GX +0.2–0.2. However, we regard the probability for chance coincidence to be not negligible so close to the Galactic center (19′). No other accurately positioned catalogued x-ray source is in the error box. The average intensities are 8 mCrab over 2–9 keV and 18 mCrab over 9–25 keV. A burst was detected coincident with the position on Mar. 21.51, with a peak intensity of ∼ 1 Crab (2–9 keV) and a duration of 15 s. GX +0.2–0.2 is not known to be a burster.”

SUPERNOVA 1998S IN NGC 3877 Visual magnitude estimates: Mar. 18.021 UT, 11.9 (B. H. Granslo,

Fjellhamar, Norway); 18.99, 12.0 (J. Shanklin, Cambridge, England); 19.93, 11.8 (J. Carvajal, Madrid, Spain); 20.93, 11.7 (Carvajal); 21.77, 12.1 (K. Hornoch, Lelekovice, Czech Republic); 22.77, 12.1 (Hornoch).

1998 March 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green